# Scalable Fine-tuning from Multiple Data Sources: A First-Order Approximation Approach

### v2

@@ Abstract
We study the problem of fine-tuning a language model (LM) for a target task by optimally using the information from $n$ auxiliary tasks. This problem has broad applications in NLP, such as targeted instruction tuning and data selection in chain-of-thought fine-tuning. The key challenge of this problem is that not all auxiliary tasks are useful to improve the performance of the target task. Thus, choosing the right subset of auxiliary tasks is crucial. Conventional subset selection methods, such as forward and backward stepwise selection, are unsuitable for LM fine-tuning because they require repeated training on subsets of auxiliary tasks. This paper introduces a new algorithm to estimate model fine-tuning performances without repeated training. Our algorithm first performs multitask training using the data of all the tasks to obtain a meta initialization. Then, we approximate the model fine-tuning loss of a subset using functional values and gradients from the meta initialization. Empirically, we find that this gradient-based approximation holds with remarkable accuracy for twelve transformer-based LMs. Thus, we can now estimate fine-tuning performances on CPUs within a few seconds. Finally, we fine-tune the pretrained base model for once on the selected subset of tasks. We conduct extensive experiments to validate this approach, delivering a speedup of 30×30$\times$30 × over conventional subset selection while incurring only $1\%$ error of the true fine-tuning performances. In downstream evaluations involving both instruction tuning and chain-of-thought fine-tuning, this loss-based selection approach improves over prior gradient or representation similarity-based methods for subset selection by up to $3.8\%$.
@@ 1 Introduction
Fine-tuning a language model (LM) has emerged as an effective approach for knowledge transfer on text data. As the scale of LMs keeps growing, efficient and scalable fine-tuning methods are in great demand. For instance, parameter-efficient fine-tuning with adapters or low-rank parameterization can significantly reduce the memory usage of fine-tuning on a large model Houlsby et al. (2019); Pfeiffer et al. (2020); Hu et al. (2021). In many applications, besides solving a target task of interest, one can also access several related data sources, which could be used for data augmentation. Examples include multilingual systems such as neural machine translation Neubig and Hu (2018), parsing Üstün et al. (2020), data selection Xie et al. (2023), and targeted instruction tuning Xia et al. (2024). A fundamental issue in these scenarios is selecting the beneficial data sources amongst all the available data sources, which can be formulated as a subset selection problem. In this paper, we develop efficient algorithms for subset selection in LM fine-tuning given $n$ data sources, which can scale up to handle large numbers of $n$.
@@
Classical subset selection methods, such as forward and backward stepwise selection, have been very effective for feature selection in regression analysis Hastie et al. (2009). However, these classical methods are unsuitable for LM fine-tuning because they require repeated training of the base LM on many subsets of tasks, which is not feasible for large $n$. More recently, data selection methods have been proposed based on selecting influential samples by comparing gradients Xia et al. (2024) or feature representation during training Ivison et al. (2023). These methods depend on the training procedure, leading to noisy outcomes for approximating true fine-tuning performances. In summary, existing subset selection methods require expensive computation or rely on noisy measurements of task-relatedness (Ruder, 2017; Vu et al., 2020, 2022; Zhang et al., 2023).
@@
This paper presents a new approach to scale up subset selection by quickly estimating model fine-tuning losses without running the fine-tuning procedure. The key idea is first to perform multitask training on the samples of all the tasks, leading to a meta initialization $\theta^{\star}$. From $\theta^{\star}$, we compute all the samples’ functional values and gradients, which can be used during inference. Second, given a subset of auxiliary tasks, we apply Taylor’s expansion on the loss and approximate the loss value with the function values and the gradients of the samples in the subset. After applying this approximation, we will estimate the fine-tuning losses directly with logistic regression without conducting any actual fine-tuning. Notice that this computation can be done on CPUs, and we use random projections to reduce the dimension of the regression down to a few hundred. Using this procedure, we can estimate the fine-tuning performance for each subset in just a few seconds, which is significantly faster than full-model fine-tuning. Importantly, the accuracy of this approach hinges on the quality of the first-order approximation. We find that across twelve transformer-based LMs, including Llama-3-8B, the approximation error of the gradient-based expansion is at the order of $10^{-5}$ to $10^{-3}$, in regions close to $\theta^{\star}$. Based on this estimation, we can apply a classical subset selection method, such as forward selection, to estimate the loss values of every subset encountered during the procedure. Lastly, we fine-tune a pretrained LM by combining the data from the selected subset of tasks. In summary, our approach has three stages. See Figure 1 for an illustration of our overall approach.
@@
We conduct extensive experiments to validate our approach using four benchmarks. We show that GradEx can accurately approximate fine-tuned model losses within 1% errors across various LMs. Meanwhile, GradEx can accelerate forward selection with 30$\times$ fewer FLOPs and 25$\times$ less GPU hours by circumventing full model fine-tuning. GradEx also speeds up random ensemble with 44$\times$ fewer FLOPs and 46$\times$ less GPU hours. We consider targeted instruction tuning and chain-of-thought fine-tuning for downstream evaluations of GradEx. Our approach matches the accuracy of classical subset selection methods with full fine-tuning while costing only 0.5% of the computation. On the ToxiGen and TruthfulQA benchmarks, GradEx outperforms existing selection methods that rely on gradient or representation measurements Xie et al. (2023); Xia et al. (2024) by 3.8% on average. On StrategyQA and CommonsenseQA datasets, the loss-based selection of GradEx improves chain-of-thought reasoning accuracy by 2.4% over gradient-based and feature-based selection methods.
@@
In conclusion, this paper introduces a new algorithm for subset selection in LM fine-tuning given multiple data sources. The algorithm can quickly estimate model fine-tuning performances without performing fine-tuning on each subset. We identify a linearization property of LMs (after meta-training), which is empirically verified on twelve transformer-based LMs. This empirical finding may be relevant for future developments of fast inference methods of LLMs. Our estimation algorithm unlocks classical subset selection methods such as forward & backward selection to LMs. The extensive experiments show that our approach can deliver significant speedups without losing downstream accuracy. In particular, our loss-based approach can outperform existing gradient-based or feature-based selection methods for instruction tuning and chain-of-thought fine-tuning. The code repository and detailed instructions for reproducing our findings are available at https://github.com/VirtuosoResearch/Scalable-finetuning.
@@ 2 Problem Formulation
Problem setup: Consider fine-tuning a language model to solve a target task. Suppose we also have access to $n$ source tasks. Because the information from different data sources may conflict with each other or hurt the target performance Wu et al. (2020), we would like to select a subset of the $n$ tasks. This often happens in transfer learning, for instance, working with low-resource neural machine translation Zoph et al. (2016). Because there are $2^{n}$ possible choices of subsets, performing the best subset selection is very costly.
In this paper, we develop a much faster approach for subset selection that is suitable for running on LMs. Concretely, given a base fine-tuning procedure on top of an LM (such as LoRA or QLoRA Dettmers et al. (2023)), let $f(S)$ be the model fine-tuning loss on the target task trained together with a given subset of auxiliary tasks $S\subseteq\{1,2,\dots,n\}$. A lower value of $f$ indicates that $S$ is more relevant to the target task. Best subset selection corresponds to finding a subset that minimizes $f(S)$.
2.1 Prior approaches for measuring $f(S)$
@@
$\diamondsuit$ Second, we also find that $f(S)$ is not monotone, in the sense that adding an extra task to $S$ may not necessarily reduce the value of $f$ (although it does more data). We pick a target task $t$ corresponding to “arrange” in Alpaca and find that $f(S)$ starts to increase after we add more than one task to $S$, which confirms our hypothesis. This suggests directly optimizing $f$ over the space of subsets remains challenging because $f$ is complex.
@@ 3 Our Approach
Table 1: We empirically find that the first-order approximation holds with very high accuracy within 0.25% relative distance to the meta initialization $\theta^{\star}$. As shown in the table below, the RRSS is at the order of $10^{-5}$-$10^{-3}$ when $X$ is between $0.05\%$ to $0.25\%$ distance of $\theta^{\star}$, where $\theta^{\star}$ includes all model weights in the initialization. We attribute this behavior to the highly overparameterized nature of LLMs. As a remark, we use LoRA as the base fine-tuning behavior; thus, the fine-tuned distance is small because of using LoRA. We report the average over $50$ random task subsets to ensure statistical significance. The reference for each LM can be found in Appendix D.
@@ 3.1 Multitask training on all tasks
Note that $f(S)$ is equal to the averaged $h_{X}(s,y)$ over a set of samples $(s,y)$. We examine first-order Taylor’s expansion of $h_{X}(s,y)$ centered at $\theta^{\star}$:
@@
Our key observation is that $\epsilon$ remains negligible for $X$ around $\theta^{\star}$. We report empirical measurements of $\epsilon$ across twelve LMs, evaluated on the GLUE benchmark (with $n=9$ tasks) for BERT and RoBERTa, and the Alpaca dataset with $n=38$ tasks for the rest of ten LMs. We use LoRA as the base fine-tuning procedure and see similar results with full fine-tuning. We compute the relative residual sum of squares (RRSS):
We first verify that the fine-tuned models on subsets of tasks stay close to the initialization. We fine-tune one LM on a random subset of tasks until the validation loss stops decreasing. We measure the relative distance from the fine-tuned model weights to the initialization, i.e., $\frac{\|X-\theta^{\star}\|_{{}_{F}}}{\|\theta^{\star}\|_{{}_{F}}}$ over 50 task subsets. We consider all model weights in the initialization as $\theta^{\star}$. For full model fine-tuning, the relative distance remains within 0.25% on average. For LoRA fine-tuning, the relative distance stays within 0.19%.
@@
3.2 Fast estimation for each task subset
@@
Denote the above as $\hat{L}_{S}(X)$, which varies by $S$. Let $\hat{\theta}_{S}\in{}^{p}$ be the solution from minimizing $\hat{L}_{S}(X)$. In practice, the dimension of $X$ can be huge; thus, we apply random projection to reduce dimension, which can provably preserve the accuracy of the regression problem through the Johnson-Lindenstrauss lemma Johnson (1984).
We sample a $p$ by $d$ (e.g., for $d=100$) Gaussian random matrix $P$ with each entry drawn from $N(0,d^{-1})$ and project the gradient as $P^{\top}h_{\theta^{\star}}(s,y)$. We insert the projected gradient as $\tilde{g}$ into ^$\hat{\ell}(X)$. Then, after solving the regression problem in dimension $d$, we map the minimizer $\hat{X}_{d}$ from dimension $d$ to $p$ using $P\hat{X}_{d}$. This step only takes a few seconds, which is extremely fast and is much faster than full fine-tuning, since it does not compute the gradient as both $b$ and $g$ are already computed after meta-training. Thus, this step will work using CPUs. Algorithm 1 summarizes our procedure.
@@
3.3 Accelerating subset selection for LMs
@@
GradEx-FS: During forward selection, we apply GradEx to every subset encountered in the selection procedure. After performing subset selection, we output the subset and use their data for augmentation.
@@
Figure 2: Illustrating the separation between correctly-labeled and noisy examples, using estimated values of $T$ in equation (1), compared to several existing measures. Our measure from $T$ (rightmost) leads to a much more clear separation of correct vs. noisy examples compared to the three measures. In particular, the second and third figures use gradient-based feature similarity, and our loss-based selection procedure leads to a more accurate separation than these measures.
3.4 Comparison with prior approaches
@@
In particular, our algorithm reduces $m$ model fine-tuning runs to a single meta-model training. As mentioned earlier, the additional estimation stage incurs very little overhead; after dimension reduction, solving each logistic regression problem takes less than 2 seconds per task subset. This estimation stage takes less than 10% of the total computation cost in our overall procedure. As for the meta-training stage, our algorithm uses comparable computational costs to existing data selection methods such as Xie et al. (2023) and Xia et al. (2024). Table 2 summarizes this comparison, including the number of forward passes required by each approach. Note that the number of forward passes is for one training step. For every method, the number of backward passes equals the forward passes. For the calculation of each method, see Appendix C for the details.
Table 2: Summary of runtime complexity between our algorithm and existing solutions for subset selection, as a function of the number of data sources $n$. We describe the constants in terms of forward passes each method takes. Here, $\alpha$ denotes the size of subsets sampled in random ensembles.
@@
Table 3: We report the relative RSS between $\hat{f}(S)$ and $f(S)$, measured on Alpaca and StrategyQA. For measuring speedup, we report the ratio of the number of FLOPs required for computing $f(S)$ divided by GradEx. The speedup remains the same across different models for the same dataset since the speed-up stems from a reduced number of forward and backward passes.

### v3

@@ Abstract
We study the problem of fine-tuning a language model (LM) for a target task by optimally using the information from $n$ auxiliary tasks. This problem has broad applications in NLP, such as targeted instruction tuning and data selection in chain-of-thought fine-tuning. The key challenge of this problem is that not all auxiliary tasks are beneficial in improving the performance of the target task. Thus, selecting the right subset of auxiliary tasks is crucial. Conventional subset selection methods, such as forward and backward stepwise selection, are unsuitable for LM fine-tuning because they require repeated training on subsets of auxiliary tasks. This paper introduces a new algorithm for estimating model fine-tuning performance without requiring repeated training. Our algorithm first performs multitask training using data from all tasks to obtain a meta initialization. Then, we approximate the model fine-tuning loss of a subset using functional values and gradients from the meta initialization. Empirically, we find that this gradient-based approximation holds with remarkable accuracy for twelve transformer-based LMs. Thus, we can now estimate fine-tuning performances on CPUs within a few seconds. Finally, we fine-tune the pretrained base model once on the selected subset of tasks. We conduct extensive experiments to validate this approach, delivering a speedup of 30×30$\times$30 × over conventional subset selection while incurring only $1\%$ error of the true fine-tuning performances. In downstream evaluations involving both instruction tuning and chain-of-thought fine-tuning, this loss-based selection approach improves over prior gradient or representation similarity-based methods for subset selection by up to $3.8\%$.
@@ 1 Introduction
Fine-tuning a language model (LM) has emerged as an effective approach for knowledge transfer with text data. As the scale of LMs continues to grow, efficient and scalable fine-tuning methods are in high demand. For instance, parameter-efficient fine-tuning with adapters or low-rank parameterization can significantly reduce the memory usage of fine-tuning on a large model Houlsby et al. (2019); Pfeiffer et al. (2020); Hu et al. (2021). In many applications, besides solving a target task of interest, one can also access several related data sources that can be used for data augmentation. Examples include multilingual systems such as neural machine translation Neubig and Hu (2018), parsing Üstün et al. (2020), data selection Xie et al. (2023), and targeted instruction tuning Xia et al. (2024). A fundamental issue in these scenarios is selecting the beneficial data sources amongst all the available data sources, which can be formulated as a subset selection problem. In this paper, we develop efficient algorithms for subset selection in LM fine-tuning given $n$ data sources, which can scale up to handle large numbers of $n$.
@@
Classical subset selection methods, such as forward and backward stepwise selection, have been very effective for feature selection in regression analysis Hastie et al. (2009). However, these classical methods are unsuitable for LM fine-tuning because they require repeated training of the base LM on many subsets of tasks, which is not feasible for large $n$. More recently, data selection methods have been proposed based on selecting influential samples by comparing gradients Xia et al. (2024) or feature representation during training Ivison et al. (2023). These methods depend on the training procedure, leading to noisy outcomes for approximating true fine-tuning performances. In summary, existing subset selection methods require expensive computation or rely on noisy measurements of task-relatedness (Vu et al., 2020, 2022; Zhang et al., 2023).
@@
This paper presents a new approach to scale up subset selection by quickly estimating model fine-tuning losses without running the fine-tuning procedure. The key idea is to first perform multitask training on samples from all tasks, leading to a meta initialization $\theta^{\star}$. From $\theta^{\star}$, we compute all the samples’ functional values and gradients, which can be used during inference. Second, given a subset of auxiliary tasks, we apply Taylor’s expansion on the loss and approximate the loss value with the function values and the gradients of the samples in the subset. After applying this approximation, we will estimate the fine-tuning losses directly with logistic regression without conducting any actual fine-tuning. Notice that this computation can be done on CPUs, and we use random projections to reduce the dimension of the regression down to a few hundred. Using this procedure, we can estimate the fine-tuning performance for each subset in just a few seconds, which is significantly faster than fine-tuning the full model. Importantly, the accuracy of this approach hinges on the quality of the first-order approximation. We find that across twelve transformer-based LMs, including Llama-3-8B, the approximation error of the gradient-based expansion is at the order of $10^{-5}$ to $10^{-3}$, in regions close to $\theta^{\star}$. Based on this estimation, we can apply a classical subset selection method, such as forward selection, to estimate the loss values of every subset encountered during the procedure. Lastly, we fine-tune a pretrained model by combining data from the selected subset of tasks. In summary, our approach has three stages. See Figure 1 for an illustration of our overall approach.
@@
We conduct extensive experiments to validate our approach using four benchmarks. We show that GradEx can accurately approximate fine-tuned model losses within 1% errors across various LMs. Meanwhile, GradEx can accelerate forward selection with 30$\times$ fewer FLOPs and 25$\times$ less GPU hours by circumventing full model fine-tuning. GradEx also speeds up random ensemble with 44$\times$ fewer FLOPs and 46$\times$ less GPU hours. We consider targeted instruction tuning and chain-of-thought fine-tuning for downstream evaluations of GradEx. Our approach matches the accuracy of classical subset selection methods with full fine-tuning while incurring only 0.5% of the computational cost. On the ToxiGen and TruthfulQA benchmarks, GradEx outperforms existing selection methods that rely on gradient or representation measurements Xie et al. (2023); Xia et al. (2024) by 3.8% on average. On StrategyQA and CommonsenseQA datasets, the loss-based selection of GradEx improves chain-of-thought reasoning accuracy by 2.4% over gradient-based and feature-based selection methods.
@@
In conclusion, this paper presents a novel algorithm for subset selection in LM fine-tuning, utilizing multiple data sources. The algorithm can quickly estimate model fine-tuning performances without performing fine-tuning on each subset. We identify a linearization property of LMs (after meta-training), which is empirically verified on twelve transformer-based LMs. This empirical finding may be relevant for future developments of fast inference methods of LLMs. Our estimation algorithm unlocks classical subset selection methods, such as forward & backward selection, to LMs. The extensive experiments show that our approach can deliver significant speedups without losing downstream accuracy. In particular, our loss-based approach can outperform existing gradient-based or feature-based selection methods for instruction tuning and chain-of-thought fine-tuning. The code repository and detailed instructions for reproducing our findings are available at https://github.com/VirtuosoResearch/Scalable-finetuning.
@@ 2 Problem Formulation
Consider fine-tuning a language model to solve a target task. Suppose we also have access to $n$ source tasks. Because the information from different data sources may conflict with each other or hurt the target performance Wu et al. (2020), we would like to select a subset of the $n$ tasks. This often happens in transfer learning, for instance, working with low-resource neural machine translation Zoph et al. (2016). Because there are $2^{n}$ possible choices of subsets, performing the best subset selection is very costly.
In this paper, we develop a significantly faster approach for subset selection that is well-suited for LLMs. Concretely, given a base fine-tuning procedure on top of an LM (such as LoRA or QLoRA Dettmers et al. (2023)), let $f(S)$ be the model fine-tuning loss on the target task trained together with a given subset of auxiliary tasks $S\subseteq\{1,2,\dots,n\}$. A lower value of $f$ indicates that $S$ is more relevant to the target task. Best subset selection corresponds to finding a subset that minimizes $f(S)$.
2.1 Task-relatedness measures
@@
$\diamondsuit$ Second, we also find that $f(S)$ is not monotone, in the sense that adding an extra task to $S$ may not necessarily reduce the value of $f$ (although it does more data). We pick a target task $t$ corresponding to “arrange” in Alpaca and find that $f(S)$ starts to increase after we add more than one task to $S$, which confirms our hypothesis. This suggests that directly optimizing $f$ over the space of subsets remains challenging because $f$ is complex.
@@ 3 Our Approach
Table 1: We empirically find that the first-order approximation holds with very high accuracy within 0.25% relative distance to the meta initialization $\theta^{\star}$. As shown in the table below, the RRSS is at the order of $10^{-5}$-$10^{-3}$ when $X$ is between $0.05\%$ to $0.25\%$ distance of $\theta^{\star}$, where $\theta^{\star}$ includes all model weights in the initialization. We attribute this behavior to the highly overparameterized nature of LLMs. As a remark, we use LoRA as the base fine-tuning behavior; thus, the fine-tuned distance is small because of using LoRA. We report the average over $50$ random task subsets to ensure statistical significance. The reference for each LM can be found in Appendix C.
@@ 3.1 Multitask training on all tasks
Note that $f(S)$ is equal to the average $h_{X}(s,y)$ over a set of samples $(s,y)$. We examine first-order Taylor’s expansion of $h_{X}(s,y)$ centered at $\theta^{\star}$:
@@
Our key observation is that $\epsilon$ remains negligible for $X$ around $\theta^{\star}$. We report empirical measurements of $\epsilon$ across twelve LMs, evaluated on the GLUE benchmark (with $n=9$ tasks) for BERT and RoBERTa, and the Alpaca dataset with $n=38$ tasks for the rest of ten LMs. We use LoRA as the base fine-tuning procedure and obtain similar results to those with full fine-tuning. We compute the relative residual sum of squares (RRSS):
We first verify that the fine-tuned models on subsets of tasks stay close to the initialization. We fine-tune one LLM on a random subset of tasks until the validation loss stops decreasing. We measure the relative distance from the fine-tuned model weights to the initialization, i.e., $\frac{\|X-\theta^{\star}\|_{{}_{F}}}{\|\theta^{\star}\|_{{}_{F}}}$ over 50 task subsets. We consider all model weights in the initialization as $\theta^{\star}$. For full model fine-tuning, the relative distance remains within 0.25% on average. For LoRA fine-tuning, the relative distance stays within 0.19%.
@@
3.2 Efficient estimation for one task subset
@@
Denote the above as $\hat{L}_{S}(X)$, which varies by $S$. Let $\hat{\theta}_{S}\in{}^{p}$ be the solution from minimizing $\hat{L}_{S}(X)$. In practice, the dimension of $X$ can be huge; thus, we apply random projection to reduce dimension, which can provably preserve the accuracy of the regression problem through the Johnson-Lindenstrauss lemma William and Lindenstrauss (1984).
We sample a $p$ by $d$ (e.g., for $d=100$) Gaussian random matrix $P$ with each entry drawn from $N(0,d^{-1})$ and project the gradient as $P^{\top}h_{\theta^{\star}}(s,y)$. We insert the projected gradient as $\tilde{g}$ into ^$\hat{\ell}(X)$. Then, after solving the regression problem in dimension $d$, we map the minimizer $\hat{X}_{d}$ from dimension $d$ to $p$ using $P\hat{X}_{d}$. This step only takes a few seconds, which is extremely fast and is much faster than full fine-tuning, since it does not compute the gradient as both $b$ and $g$ are already computed after meta-training. Thus, this step can be performed using CPUs. Algorithm 1 summarizes our procedure.
@@
3.3 Accelerating subset selection on LLMs
@@
GradEx-FS: During forward selection, we apply GradEx to every subset encountered in the selection procedure. After performing subset selection, we output the subset and use its data for augmentation.
@@
Figure 2: Illustrating the separation between correctly-labeled and noisy examples, using estimated values of $T$ in equation (1), compared to several existing measures. Our measure from $T$ (rightmost) leads to a much clearer separation of correct vs. noisy examples compared to the three measures. In particular, the second and third figures utilize gradient-based feature similarity, and our loss-based selection procedure yields a more accurate separation than these measures.
3.4 Comparison with prior methods
@@
In particular, our algorithm reduces $m$ model fine-tuning runs to a single meta-model training. As mentioned earlier, the additional estimation stage incurs very little overhead; after dimension reduction, solving each logistic regression problem takes less than 2 seconds per task subset. This estimation stage takes less than 10% of the total computation cost in our overall procedure. Regarding the meta-training stage, our algorithm incurs comparable computational costs to existing data selection methods, such as those presented by Xie et al. (2023) and Xia et al. (2024). Table 2 summarizes this comparison, including the number of forward passes required by each approach. Note that the number of forward passes is for one training step. For every method, the number of backward passes equals the forward passes. For the calculation of each method, see Appendix B for the details.
Table 2: Summary of runtime complexity between our algorithm and existing solutions for subset selection, as a function of the number of data sources $n$. We describe the constants in terms of the forward passes each method takes. Here, $\alpha$ denotes the size of subsets sampled in random ensembles.
@@
Table 3: We report the relative RSS between $\hat{f}(S)$ and $f(S)$, measured on Alpaca and StrategyQA. For measuring speedup, we report the ratio of the number of FLOPs required for computing $f(S)$ to the number of FLOPs required by GradEx. The speedup remains the same across different models for the same dataset, as it stems from a reduced number of forward and backward passes.